OpenStudy (anonymous):

what is the solution of an equation -7+3b=4b-8

OpenStudy (bibby):

\[-7+3b=4b-8\] \[+7-3b=-3b+7\]

OpenStudy (anonymous):

b=1

OpenStudy (bibby):

prove it

OpenStudy (anonymous):

-7+3(1)=4(1)-8 both sides =-4

OpenStudy (anonymous):

solve for b